hello, i am looking for an answer to the question in maths which is related to the chapter pair of staight lines. the question is the slope of one line of 3x square + 2hxy + 8y square is double the slope of the other line then h square =

hello,
           i am looking for an answer to the question in maths which is related to the chapter pair of staight lines. the question is
 
the slope of one line of 3x square + 2hxy + 8y square is double the slope of the other line then h square =

Grade:9

2 Answers

Nayak Roshan
7 years ago
He let the slopes of two lines be m1 and m2. Therefore given is 2m1=m2.
In the given equation there are two straight lines. Buy some derivation you get m1 + m2 = -2h/b and m1m2=a/b. 
Here the equation is of the form ax+ 2hxy + by2.  which gives a=3,b=8.
Solving: m1+m2 = -2h/8 , but 2m1=m2.
Therefore  3m1= -2h/8
m1=-2h/24. (Eq-1).
And m1m2=a/b,
Therefore m1 (2m1)=3/8
2ml2=3/8
m1= root3/4. (Eq-2).
Equating 1 and 2 , -2h/24=root3/4
-2h=6 (root3)
h= (-3(root3))
h2=(-3(root3))2
h2=9×3=27
For the proof related to those formulas please refer your notes or coaching material but I have solved you the problem....
HOPE THIS HELPED YOU....
Nayak Roshan
7 years ago
Are you the candidate of BASE BANGALORE. And if yes please refer to that material it's good even i am a candidate of that coaching institute.....